# Prospects for sub-EW ALP searches via $\gamma+b\bar{b}$ signatures at the LHC using jet substructure techniques

### v1

@@ 4.3 Interpretation of the results
In order to assess the potential impact of these limits, it is interesting to visualise them along with existing constraints on the ALP parameter space. These existing constraints include limits from astrophysics and colliders which are summarised in Fig. 10, taken from Ref. Bharucha:2022lty .555 We are very grateful to Sophie Mutzel for providing the data required to reproduce these constraints. More precisely, the $m_{a}$ $g_{\rm aff}$ parameter space is constrained by: astrophysical constraints from SN1987A Kelly:2020dda ; Ertas:2020xcc and Horizontal Branch (HB) stars Pospelov:2008jk ; Avignone:1986vm , excluding regions below $m_{a}\lesssim$ 100 MeV and $\sim$100 keV, respectively; collider constrains excluding regions above $g_{\rm aff}\sim 10^{-5}$ GeV-1 for ALP masses below 10 GeV, which include electron beam dump experiments from SLAC E137 Bjorken:1988as and heavy meson decays coming from LHCb, CHARM, NA62, NA48 LHCb:2016awg ; LHCb:2015nkv ; Dobrich:2018jyi ; NA62:2021zjw ; NA482:2009pfe ; NA482:2016sfh . Importantly, these bounds only apply below $m_{a}=10$ GeV, whereas our derived projected limits constrain the ALP mass range between 12 GeV and 100 GeV. In Fig. 10, the dashed-dot blue line corresponds to the upper limit at $\sqrt{s}=13$ TeV with $\mathcal{L}=300$ fb-1. While the dashed blue and red lines represent projections for $p_{\rm T,\gamma}>160$ GeV and 100 GeV, respectively, at $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1. Projections for the LHC at $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1 show that searches for the $b\bar{b}\gamma$ final state can probe the unexplored ALP mass regions ($\mathcal{O}(10)$ GeV) with a maximum reach of $g_{\rm aff}\sim 0.026$ GeV-1 under improved trigger requirements, $p_{\rm T,\gamma}>100$ GeV.
@@
We note that since Ref. Bharucha:2022lty appeared, direct constraints on ALP-fermion couplings have been further investigated. In Ref. Esser:2023fdo the authors studied a scenario in which the only non-zero coupling of the ALP to SM fermions was that to top quarks, i.e $g_{\rm att}\neq 0$ only. This has the important consequence that the ALP is long-lived with respect to the scale of the LHC detectors. As this is not the case in our model, the derived direct bounds do not apply here. Note that indirect bounds from the LHC on the ALP-top coupling were also derived in Ref. Esser:2023fdo , and from low-energy, Higgs and top data to constrain the ALP-fermion couplings in Ref. Biekotter:2023mpd . In the former, indirect effects of the ALP on top pair production and gauge boson pair production are used to obtain limits on the ALP-top coupling. In the latter, bounds on the ALP couplings to up-type quarks ($C_{u}$), down-type quarks ($C_{d}$) and charged leptons ($C_{e}$) were obtained by taking into account renormalization group flow of these couplings into the SM effective field theory (SMEFT) Wilson coefficients, on which the constraints were already known. It turns out that particularly strong bounds on $C_{u}/f_{a}$ were obtained, stemming from the two-loop contribution of the ALP-top coupling to the $W$ mass. We note that not only do our direct constraints provide complementary information to these indirect bounds, they in addition provide the first direct constraints on the coupling of the ALP to down-type quarks in this region of $m_{a}$.
Figure 10: Constraints from astrophysical and collider experiments Bharucha:2022lty together with our derived bounds on the ALP-fermion coupling, $g_{\rm aff}$ as a function of the ALP mass, $m_{a}$. The LHC bounds for $p_{\rm T,\gamma}>160$ GeV is shown with dashed-dotted and dashed blue color for $\sqrt{s}=13$ TeV with $\mathcal{L}=300$ fb-1 and $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1, respectively. The improvement with $p_{\rm T,\gamma}>100$ GeV requirement is shown with red dashed line at $\sqrt{s}=14$ TeV with $\mathcal{L}=3000$ fb-1.
@@ 5 Summary and outlook
We have studied the LHC prospects for an ALP, a light pseudoscalar which primarily decays to a $b\bar{b}$ final state, in the mass range between 10 and 100 GeV. We considered a simple DFSZ-inspired scenario where the ALP only couples to SM fermions and this coupling is proportional to the fermion Yukawa.
@@
A crucial aspect in our analysis involves the $b$-tagging of the AK1.0 jet. We performed it by constructing at least two variable radius subjets inside the AK1.0 jet and matched them to a hadron in the event, resulting in 6 possible di-flavour combinations for the multijet background: $bb$, $bc$, $bl$, $cc$, $cl$, and $ll$. Standard tagging efficiencies for single-jet flavour can not be applied here. We included the corresponding tagging efficiencies for the signal and multijet rejection to those di-flavour combinations for the background provided by the ATLAS collaboration ATL-PHYS-PUB-2023-021 .
The final signal efficiency and background yields are calculated by performing a Crystal Ball fit to the signal ${\mathrm{m}}_{\mathrm{SD}}$ distribution, and selecting the $1\sigma$ window around the fitted mean value. The upper limits are set at $95\%$ confidence level on the $b\bar{b}\gamma$ production cross-section, $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$. The HL-LHC limits vary between 185 pb and $4\cdot 10^{3}$ pb, without systematics, 440 (1014) pb and $8.7\cdot 10^{3}$ ($1.9\cdot 10^{4}$) pb, with $2\%$ ($5\%$) systematics, for ALP masses between 20 GeV and 70 GeV. Translating the cross-section limits to the ALP-fermion coupling, the upper bound on $g_{\mathrm{aff}}$, in the ALP mass range between 20 GeV and 70 GeV, is (0.053(0.053( 0.053 - $0.069$) GeV-1 at $95\%$ CL. The $g_{\mathrm{aff}}$ upper limits worsen upon adding systematic uncertainty, and the maximum obtainable reach is 0.076 (0.11) GeV-1, around $m_{a}\sim$ 40 GeV, upon adding $2\%$ ($5\%$) systematics.
Comparing these limits to the existing LHC searches at the current run, we find that our analysis provides significantly better exclusion limits on the $b\bar{b}\gamma$ production cross-section. In addition, we complement our analysis with a reduced $p_{\mathrm{T},\gamma}$ threshold, to assess the capability of our analysis in increasing the sensitivity to low mass scenarios. Indeed, our study demonstrates that the sensitivity to the cross-section improves by a factor of four at $m_{a}=12$ GeV. The coupling $g_{\mathrm{aff}}$ can be excluded above $0.026$ GeV-1 at $95\%$ CL for $m_{a}\sim 30-40$ GeV. Thus, LHC searches on the proposed $b\bar{b}\gamma$ final state have sensitivity to the considered mass range which has not been explored yet by other complementary collider searches.

### v3

@@ 4.3 Interpretation of the results
In order to assess the potential impact of these limits, it is interesting to visualise them along with existing constraints on the ALP parameter space. These existing constraints include limits from astrophysics, beam-dump and flavour experiments for the region $m_{a}\lesssim 5$ GeV, as calculated in Ref. Bharucha:2022lty and ATLAS/CMS bounds for larger ALP masses, calculated here. More precisely, the $m_{a}$ - $g_{\rm aff}$ parameter space is constrained by: astrophysical constraints from SN1987A Kelly:2020dda ; Ertas:2020xcc and Horizontal Branch (HB) stars Pospelov:2008jk ; Avignone:1986vm ; electron beam dump experiments from SLAC E137 Bjorken:1988as and heavy meson decays coming from LHCb, CHARM, NA62, NA48 LHCb:2016awg ; LHCb:2015nkv ; Dobrich:2018jyi ; NA62:2021zjw ; NA482:2009pfe ; NA482:2016sfh . For ALP masses $\gtrsim 5$ GeV , the bounds are derived from BSM searches at LHC, including dijet searches CMS:2019xai ; CMS:2017dcz , di-photon searches Mariotti:2017vtv ; ATLAS:2022abz , top-quark pair production with a lepton pair ($t\bar{t}\ell^{+}\ell^{-}$CMS:2024ulc , and the $h\to$ BSM search ATLAS:2022vkf . In addition, SM measurements of $t\bar{t}$ production in association with a bottom pair ($t\bar{t}b\bar{b}$CMS:2020utv , four-top-quark production ($t\bar{t}t\bar{t}$AlvarezGonzalez:2024tot ; CMS:2023zdh , and $t\bar{t}$ production ATLAS:2019hxz are used to constrain part of the model parameter space.
Figure 10: Constraints from astrophysical and collider experiments on the ALP-fermion coupling, $g_{\rm aff}$ as a function of the ALP mass, $m_{a}$. The astrophysical constraints include SN1987A Kelly:2020dda ; Ertas:2020xcc , Horizontal Branch (HB) stars Pospelov:2008jk ; Avignone:1986vm . The collider constrains come from electron beam dump experiments from SLAC E137 Bjorken:1988as , heavy meson decays coming from LHCb, CHARM, NA62, NA48 LHCb:2016awg ; LHCb:2015nkv ; Dobrich:2018jyi ; NA62:2021zjw ; NA482:2009pfe ; NA482:2016sfh , di-photon searches Mariotti:2017vtv ; ATLAS:2022abz , di-jet searches CMS:2019xai ; CMS:2017dcz , $h\to$ BSM ATLAS:2022vkf , $t\bar{t}b\bar{b}$ CMS:2020utv and $t\bar{t}t\bar{t}$ AlvarezGonzalez:2024tot ; CMS:2023zdh production.
In Fig. 10, we summarise the astrophysics and collider constraints that are adapted  [^5] We are very grateful to Sophie Mutzel for providing the data required to reproduce these constraints. from Ref. Bharucha:2022lty   along with our derived collider bounds for larger ALP masses in the $g_{\rm aff}-m_{a}$ plane. The astrophysical bounds from SN1987A and HB stars exclude regions below $m_{a}\lesssim$ 100 MeV and $\sim$100 keV, respectively. The flavor constrains for $m_{a}\lesssim 5$ GeV, exclude regions above $g_{\rm aff}\sim 10^{-5}$ GeV-1. Above the ALP mass of $\sim 5$ GeV, we include the following collider constraints:
  • The $\mathbf{h\to}$ BSM bound arises as the SM Higgs can decay to ALPs via fermion loops and thus constrain the ALP-fermion coupling. We first calculate the $h\to aa$ branching ratio, considering the top- and bottom-quark loop contributions. We use FeynRules Alloul:2013bka to generate the model file for FeynArts Hahn:2000kx , then FeynArts, FormCalc and LoopTools Hahn:1998yk packages to generate and calculate the Feynman diagrams. We take a similar approach as in Ref. Blasi:2023hvb : we evaluate the scale-dependent loop diagrams, taken for the renormalisation scale $\mu=m_{h}$. An upper limit on the branching ratio $h\to$ BSM, Br($h\to$ BSM)<0.073<0.073, was derived in Ref. Blasi:2023hvb from ATLAS measurements ATLAS:2022vkf . This is effectively an upper bound on the derived Br($h\to aa$), as here this is the dominant BSM decay channel of the Higgs. We show this bound in Fig. 10 in the $g_{\rm aff}-m_{a}$ plane. Notice however that tree-level contributions to $h\to aa$ are also possible due to the dim-6 operator $(\partial_{\mu}a)(\partial_{\mu}a)H^{\dagger}H$, where the corresponding Wilson coefficient is independent of $g_{\rm aff}$. Therefore the $h\to$ BSM constraint can be relaxed by a cancellation between these two independent contributions and we hence show it by a dotted-shaded line in Fig. 10.
  • To constrain $g_{\rm aff}$ from the $\mathbf{t\bar{t}b\bar{b}}$ and $\mathbf{t\bar{t}t\bar{t}}$ processes, we calculate the production cross-section in our ALP model using MadGraph and demand that they agree with the experimentally measured cross-section. In the case of the $t\bar{t}b\bar{b}$ process, the full phase space cross-section as defined in Ref. CMS:2020utv is evaluated with the following requirements on the jets: $p_{\rm T}>$ 20 GeV and $|\eta|<$ 2.4. An upper limit on $g_{\rm aff}$ is set by comparing the obtained cross-section to the measured value of 4.54 $\pm$ 0.35 (stat.) $\pm$ 0.56 (syst.) $\pm$ 0.66 (MG5) pb CMS:2020utv , as illustrated in Fig. 10 by the magenta region. As the $t\bar{t}\ell^{+}\ell^{-}$ are similar to those from $t\bar{t}b\bar{b}$ (see Ref. Blasi:2023hvb ), we only calculate the former. In the case of $t\bar{t}t\bar{t}$ production, the measured cross-section is 17 $\pm$ 4 (stat.) $\pm$ 3 (syst.) fb AlvarezGonzalez:2024tot ; CMS:2023zdh . The corresponding upper bounds on $g_{\rm aff}$ from the $t\bar{t}t\bar{t}$ is shown in Fig. 10 by the yellow region.
  • The dijet resonance searches from CMS CMS:2019xai ; CMS:2017dcz are shown by the blue and green regions in Fig. 10, similar to the left panel of Fig. 8. In both ATLAS and CMS analyses ATLAS:2022abz ; Mariotti:2017vtv , the di-photon constraints are derived assuming a gluon fusion production mechanism, whereas in the ALP model considered here, the ALP production is initiated from quark annihilation. Thus, we reinterpret these limits by scaling the production cross-section times branching ratio, with the help of Ref. Blasi:2023hvb , and derive the bound in the $g_{\rm aff}-m_{a}$ plane, shown by the light-blue and red regions.
Note that $t\bar{t}$ production could also exclude regions in the parameter space considered. However, we do not derive the bound here as the ALP contribution does not arise at tree level, necessitating a calculation of the interference between the SM and BSM loop contribution Maltoni:2024wyh to reinterpret the bound in our model, which is beyond the scope of the present work.
Figure 11: The collider constraints on the ALP-fermion coupling, $g_{\rm aff}$ as a function of the ALP mass, $m_{a}$, in the mass range between 1210012-100 GeV, is compared with our derived bounds. The collider constraints are the same as described in Fig. 10, while the derived bounds are shown with blue and red lines (as discussed in Fig. 9). On the left panel, the bounds take into account the ALP couplings to top and bottom quark, while the ALP only couples to bottom quark for the constraints shown on the right panel.
While the ALP-top quark coupling has little relevance for our signal, it is crucial for several of the collider bounds mentioned above. In Fig. 11, we therefore focus on the region $10\lesssim m_{a}\lesssim 100$ GeV for the cases $g_{\rm att}=g_{\rm abb}$ (left panel) and $g_{\rm att}=0$ (right panel), in order to compare the existing constraints with our derived projected limits, shown by the blue and red dashed lines for $p_{\rm T,\gamma}>160$ GeV and $p_{\rm T,\gamma}>100$ GeV, respectively. The case of an ALP coupling to both the up-type and down-type quarks ($g_{\rm att}=g_{\rm abb}$), corresponding to Fig. 10, is illustrated in the left panel of Fig. 11. We note that the constraint from $t\bar{t}b\bar{b}$ excludes a larger region than that from $\gamma+b\bar{b}$ derived in this work. On the other hand, the $t\bar{t}t\bar{t}$ bound is flat over the considered mass range and weaker than the $\gamma+b\bar{b}$ sensitivity for $p_{\rm T,\gamma}>100$ GeV, for ALP masses in the range 20-60 GeV. Below $m_{a}\sim 65$ GeV, the di-photon constraint is weaker than the $\gamma+b\bar{b}$ bounds.
The right panel in Fig. 11 summarises the constraints on a class of models where the ALPs interact dominantly with down-type quarks ($g_{\rm att}=0$). As expected, all the previously discussed collider constraints become comparatively weaker than the $\gamma+b\bar{b}$ projected sensitivity for HL-LHC. In particular, for the $t\bar{t}b\bar{b}$ process, once we set the ALP-top coupling to zero, the only contributions are from quark-initiated diagrams, which involve top-quark emission from SM gauge or Higgs bosons. In the case of $h\to aa$, the bound only includes the bottom quark loop contribution. Similarly, the process $a\to\gamma\gamma$ is induced only by the bottom quark loop. It should also be noted that when the ALP-top coupling is zero, the ALP contribution to the $t\bar{t}t\bar{t}$ production vanishes. In contrast, our signal is only weakly affected when the ALP coupling to up-type quarks is switched off: the production cross-section is reduced roughly by a factor two, corresponding to the up-type quark content of the proton, and this relaxes the upper bound on $g_{\rm abb}$ by roughly a factor $\sqrt{2}$.
@@
Other constraints on the ALP-top coupling have been investigated since Ref. Bharucha:2022lty appeared. In Ref. Esser:2023fdo the authors studied a scenario in which the only non-zero coupling of the ALP to SM fermions was that to top quarks, i.e $g_{\rm att}\neq 0$ only. This has the important consequence that the ALP is long-lived with respect to the scale of the LHC detectors. As this is not the case in our model, the derived direct bounds do not apply here. Indirect bounds from the LHC on the ALP-top coupling were also derived in Ref. Esser:2023fdo , and from low-energy, Higgs and top data on the ALP-fermion couplings in Ref. Biekotter:2023mpd . In the former, indirect effects of the ALP on top pair production and gauge-boson pair production are used to obtain limits on the ALP-top coupling. In the latter, bounds on the ALP couplings to up-type and down-type quarks as well as charged leptons were obtained by taking into account renormalization group flow of these couplings into the SM effective-field-theory (SMEFT) Wilson coefficients, on which the constraints were already known. It turns out that particularly strong bounds on couplings of ALPs to up-type were obtained, stemming from the two-loop contribution of the ALP-top coupling to the $W$ mass. While these bounds are clearly interesting, reproducing them for our model is beyond the scope of our work. We remark that the direct searches we consider provide important complementary information to these indirect bounds.
@@ 5 Summary and outlook
We have studied the LHC prospects for an ALP, a light pseudoscalar which primarily decays to the $b\bar{b}$ final state, in the mass range 10 to 100 GeV. We considered a simple DFSZ-inspired scenario where the ALP only couples to SM fermions and this coupling is proportional to the fermion Yukawa.
@@
A crucial aspect in our analysis involves the $b$-tagging of the AK1.0 jet. We performed it by constructing at least two variable radius subjets inside the AK1.0 jet and matched them to a hadron in the event, resulting in 6 possible di-flavour combinations for the multijet background: $bb$, bcbc, blbl, cccc, clcl, and $ll$. Standard tagging efficiencies for single-jet flavour can not be applied here. We included the corresponding tagging efficiencies for the signal and multijet rejection to those di-flavour combinations for the background provided by the ATLAS collaboration ATL-PHYS-PUB-2023-021 .
The final signal efficiency and background yields are calculated by performing a Crystal Ball fit to the signal ${\mathrm{m}}_{\mathrm{SD}}$ distribution, and selecting the $1\sigma$ window around the fitted mean value. The upper limits are set at $95\%$ confidence level on the $b\bar{b}\gamma$ production cross-section, $\sigma(pp\to a\gamma\to b\bar{b}\gamma)$. The HL-LHC limits vary between 185 pb and $4\cdot 10^{3}$ pb, without systematics, 440 (1014) pb and $8.7\cdot 10^{3}$ ($1.9\cdot 10^{4}$) pb, with $2\%$ ($5\%$) systematics, for ALP masses between 20 GeV and 70 GeV. Translating the cross-section limits to the ALP-fermion coupling, the upper bound on $g_{\mathrm{aff}}$, in the ALP mass range between 20 GeV and 70 GeV, is $(0.065$ - $0.047$) GeV-1 at $95\%$ CL. The $g_{\mathrm{aff}}$ upper limits worsen upon adding systematic uncertainty, and the maximum obtainable reach is 0.066 (0.098) GeV-1, around $m_{a}\sim$ 40 GeV, upon adding $2\%$ ($5\%$) systematics.
Comparing these limits to the existing LHC searches at the current run, we find that our analysis provides significantly better exclusion limits on the $b\bar{b}\gamma$ production cross-section. In addition, we complement our analysis with a reduced $p_{\mathrm{T},\gamma}$ threshold, to assess the capability of our analysis in increasing the sensitivity to low mass scenarios. Indeed, our study demonstrates that the sensitivity to the cross-section improves by a factor of four at $m_{a}=12$ GeV. The coupling $g_{\mathrm{aff}}$ can be excluded above $0.023$ GeV-1 at $95\%$ CL for $m_{a}\sim 30-40$ GeV.
We compare the results of this analysis to complementary existing astrophysical, flavor and collider searches. While the astrophysical and flavor bounds dominate around $m_{a}\lesssim m_{b}$, collider searches from the ATLAS and CMS experiment become important above an ALP mass of $m_{a}\gtrsim 10$ GeV. Comparing existing ATLAS and CMS constraints to our derived limits, the current bounds from $t\bar{t}b\bar{b}$ searches is slightly stronger than our bound. However, these collider bounds heavily rely on the coupling of the ALP to the top quark. We find that, in models with no ALP coupling to up-type quarks, the analysis we propose would be the most powerful direct search for the ALP.